6. Solve the differential equation by variation of parameters. y – 2y + y = 1+x2
Answer #1

Variation of parameters method Consider second order differential equation with constant coefficients as: day tale + Qoy = Q(day - 2 dy dax ex x dacz +y = H22 9 H22 for the complementary furetim Yc: Auxiliary equation of the differential equation wilxe22 1 wi It x² 22 o Also Wat ex y, 0 y Q (0) ས་ ex Hx2 14x² -> W21 2x е H22 27 So, 41 = Will do af te ha dx da → 4 = 6 - 2 uTherefore, Yp = 4, 4 tylly, Y = et Y₂ = xet » Yp=e* (- log (143) + xe tanse Ye = Getxe or Yc = (a+Gx) et Hence, required solu